OpenStudy (anonymous):

let f(x)=4x+1 and g(x)=4+x^2. Compute (g o f)(x) and f(g(-2)). Please show all work so I understand how you got your answer. Thanks so much !

hero (hero):

(g o f)(x) = g(f(x)) = 4+ (4x+1)^2 = 4 + 16x^2 + 8x + 1 = 16x^2 + 8x + 5

OpenStudy (anonymous):

I found you Hero, you know how to play hide n' seek, but I am a very good seeker.

hero (hero):

What the? lol

OpenStudy (anonymous):

xD

hero (hero):

16(-2)^2 + 8(-2) + 5 = 64 + 5 - 16 = 48 +5 = 53

OpenStudy (anonymous):

Can you answer this quote? "Don't be afraid to give up the good to go for the great" - John D. Rockefeller

hero (hero):

It means that if you find a woman you know is great (dynamic), don't pass up on her because otherwise, you'll be left with only the good (average) ones.

OpenStudy (anonymous):

Ohh okay.